HIGH RESOLUTION INFRARED SPECTRA OF AR-WATER AND NE-WATER AT 6 μm

\underline{X. LIU} and Y. XU, Department of Chemistry, University of Alberta, Edmonton, AB T6G 2G2.

Ar- and Ne-water are highly floppy van der Waals complexes where the water subunit experiences nearly free internal rotation. Their ro-vibrational energy levels are characterized by the internal rotor states of the water subunit within the complex and a pseudo-diatomic rotational energy Hamiltonian. Large amplitude motions of the complexes lead to strong perturbations, such as Coriolis coupling and angular-radial coupling among the internal rotor states and the van der Waals bending and stretching states. Mid-infrared spectra of Ar- and Ne-water were measured with a direct absorption spectrometer with an external cavity quantum cascade laser at 6 μm and a 366-pass astigmatic absorption cell.^{a} The scan-to-scan frequency instability of the laser was addressed with a “on-the-fly” calibration procedure. The infrared spectrum of Ar-water has been studied by Weida and Nesbitt,^{b} in which the Σ1_{10} and Π1_{10} states have been identified. At least three new overlapping bands at 1630 cm^{−1} have been observed and two of them have been tentatively assigned to the n = 1, Σ1_{01} ← Π1_{10} and Σ1_{10} ← Σ1_{01} bands. The n = 1, Π1_{01} ← Σ1_{01} band that was missing in the previous study was found at 1639 cm^{−1}. Four new bands in the 1645-1665 cm^{−1} region have been observed and assigned to the Π2_{12} ← Π1_{01}, Σ2_{12} ← Σ1_{01}, Π2_{12} ← Σ1_{01}, and n = 1, Σ1_{11} ← Σ0_{00}. A global fit of the microwave, far-infrared, nearinfrared and mid-infrared data was performed with Pickett’s SPFIT program to determine the spectroscopic constants of these levels. Infrared spectrum of Ne-water is analogous to that of Ar-water. The Ne-water PES^{c} is much shallower than that Ar-water. As a result, there are fewer number of internal rotor states supported by the surface. Indeed, only the Π1_{10} ← Σ1_{01}, Σ1_{10} ← Π1_{01}, n = 1, Σ0_{00} ← Σ0_{00}, and Π1_{11} ← Σ0_{00} bands were observed in our measurements. Severe spectral perturbations have been detected and analyzed. No infrared transitions of He-water were detected in the same frequency region. This is consistent with the previous theoretical calculation which predicted that the He-water potential supports only one bound state and no infrared transitions could be observed in this region because of the b− type selection rule.

aX. Liu, Y. Xu, Z. S., W. S. Tam, I. Leonov, Appl. Phys. B, \underline{102}, 629, 2011

bM. Weida, D. Nesbitt, J. Chem. Phys. \underline{106}, 3078, 1997

cA. Haskopoulos, J. Phys. Chem. A, \underline{114}, 8730, 2010